El pH de una solución de 0.02 M de un ácido débil desconocido es 3.7. ¿cuál es el pKa de este ácido?


1
Expert's answer
2022-01-19T10:11:59-0500

Solution

  • [H+] = "10^{-pH}" so since your acid was measured at 3.7
  • you place that into your equation for the pH
  • "10^{-3.7}" placing that into your calculator you will get "19.95\times10^{-5}" .
  • Now "Ka =\frac{ [H+][A-]}{[HA]}"
  • So we get as "\frac{(19.95\times10^{-5}M)^2}{0.02M}" after calculating you will get "9.95*10^{-5}"
  • "p_{ka}=-log_{10}(19.95\times10^{-5})\\ = 3.7"
  • Suppose,
  • "[H+]=10^{-3.7}\\=19.95*10^{-5 }M"
  • C*α =[H+]
  •  
  • "So,\u03b1 =9.83*10^{-3}"
  •  
  • "Ka = C*α^2/(1-α) =3.1587*10^{-8
 }
\\pKa =7.50048"